Budgeting and Variance Analysis

In an effort to better plan for and control OR costs, SHH management asked Jack to calculate the flexible budget variance (i.e., flexible budget costs - actual costs) for OR nursing costs, including the price variance and efficiency variance. Given that Jack is interested in comparing the reported costs of both systems, he decided to prepare the requested OR variance analysis for both the current cost system and the vital-signs costing system. In addition, Jack chose to use each cost system’s estimate of the cost per OR nursing hour as the standard cost per OR nursing hour. Jack collected the following additional information for use in preparing the flexible budget variance for both systems:

Actual number of surgeries performed = 950
Standard number of nursing hours allowed for each OR surgery = 5
Actual number of OR nursing hours used = 5,000
Actual OR nursing costs = $190,000

Enter a favorable variance as a negative amount, and an unfavorable variance as a positive amount. If there is no variance, enter "0" and select "No variance" from the dropdown.

7. For the OR service line, use the information above and the cost per OR nursing hour under the current cost system to calculate the

a. flexible budget variance. (Hint: Use your answer to Requirement 3 as the standard cost per OR nursing hour for the current cost system.)
$  

b. price variance.
$  

c. efficiency variance.
$  

8. For the OR service line, use the information above and the cost per OR nursing hour under the vital-signs costing system to calculate the

a. flexible budget variance. (Hint: Use your answer to Requirement 6 as the standard cost per OR nursing hour for the vital signs cost system.)
$  

b. price variance.
$  

c. efficiency variance.
$  

Discussion of Reported Costs and Variances from the Two Systems

9. Consider SHH’s need to control its skyrocketing costs, Jack’s discussion with experienced nurses regarding their use of hospital resources, and the reported costs that you calculated from each cost system. Based on these considerations, which cost system (current or vitalsigns) should Jack choose? Briefly explain the reasoning behind your choice.

a. (choose 1.vital signs or 2. current) costing system should more accurately allocate costs to service lines because its cost allocation base.

b. (choose 1.both, 2.current, or 3 vital signs) uses only one cost driver and   cost effective.

c. The more accurate (choose 1. vital signs, or 2. current) system should generate a more accurate estimate of the cost per nursing hour, which affects the budgeting process, because the portion of costs allocated to each service line, ER.

10. What does each of the calculated variances suggest to Jack regarding actions that he should or should not take with respect to investigating and improving each variance? Also, briefly explain why the variances differ between the two cost systems.

a. The overall current system’s OR flexible budget variance ($47,500) is very- (1.large and unfavorable, 2.large and favorable, 3.small and favorable or 4. zero) , suggesting that the subvariances (price variance and efficiency variance) should be calculated.

b. The current system’s OR price variance ($40,000) is very large and unfavorable, suggesting that the nursing hiring manager negotiated a-(1.bad or 2.good) price and that nursing hour pay cuts might be necessary.

c. The current system’s OR efficiency variance ($7,500) is-(Choose 1. moderate and unfavorable, 2. large and favorable, 3 large and unfavorable 4. zero) , suggesting that the operating room manager used too many OR nursing hours for the actual number of surgeries performed.

d. The overall vital signs’ OR flexible budget variance is(Choose 1. moderate and unfavorable, 2. large and favorable, 3 large and unfavorable 4. zero)  , and suggests that nothing needs to be investigated further.

e. The vital signs’ OR price variance ($10,000) is (1.large and unfavorable, 2.large and favorable, 3.small and favorable  large and favorable, suggesting that the nursing hiring manager negotiated a good price.

f. The vital signs’ OR efficiency variance ($10,000) is(1.large and unfavorable, 2.large and favorable, 3.small and favorable or 4. zero)  , suggesting that the operating room manager used too many OR nursing hours for the actual number of surgeries performed. In addition, it would be unwise had Jack decided to end the variance analysis after seeing that the flexible budget variance was zero. Only after continuing on with the analysis to calculate the price and efficiency variances would Jack realize that the zero flexible budget variance was the result of two large offsetting variances, both of which likely require further investigation and attention.

g. Overall, the two cost systems produce -(Choose 1.different or 2.same) reported costs of the two service lines, ER and OR. The current system assigns nursing costs equally because the ER and OR have the same number of patients. Alternately, the vital-signs system assigns (Choose 1. twice or 2 thrice) as much of the nursing costs to the OR because the OR requires  (Choose 1. twice or 2 thrice) as many vital signs checks of its patients as the ER does of its patients. In addition, the two systems produce (Choose 1. different or 2. same) estimates of the cost incurred by the hospital per OR nursing hour. When used as the standard costs in the budgeting process, these (Choose 1. different or 2. same) reported costs, lead to very different flexible budget variances and price and efficiency variances for the OR service line. Therefore, the managerial accountant should be very careful when constructing a cost system and be sure that the chosen allocation bases are as accurate as possible to match the underlying resource consumption patterns of the business environment. Choosing different cost allocation bases usually will result in differences in reported service line costliness and various variances, which can have ramifications for numerous managers (e.g., purchasing managers responsible for price variances, production managers responsible for efficiency variances, other managers responsible for making service line mix decisions, etc.)
Answer #1

Given that the following information

In an effort to better plan for and control OR costs, SHH management asked Jack to calculate the flexible budget variance (i.e., flexible budget costs - actual costs) for OR nursing costs, including the price variance and efficiency variance. Given that Jack is interested in comparing the reported costs of both systems, he decided to prepare the requested OR variance analysis for both the current cost system and the vital-signs costing system.

7. Flexible budget Variance:

Actual cost = 190,000

Budgeted cost = Budgeted hours * Standard cost

Budgeted hours = Budgeted no of Surgeries * hours required per surgery = 950*5 = 4,750

Budgeted cost = 4,750*30 = 142,500

Variance = 142,500-190,000 = 47,500 Unfavourable

Price variance = Actual hours * (Actual rate - Standard rate) = 5,000 * (38-30) = 40,000 U

Actual rate = 190,000/ 5,000 = 38

Efficiency variance = Standard rate * (Actual hours - Standard hours) = 30* (5,000 - 4,750) = 7,500 U

8. Variances under vital signs check cost method:

Flexible budget Variance

Actual cost = 190,000

Budgeted cost = Budgeted hours * Standard cost

Budgeted hours = Budgeted no of Surgeries * hours required per surgery = 950*5 = 4,750

Budgeted cost = 4,750*40 = 190,000

Variance = 190,000-190,000 = 0

Price variance = Actual hours * (Actual rate - Standard rate) = 5,000 * (38-40) = 10,000 F

Actual rate = 190,000/ 5,000 = 38

Efficiency variance = Standard rate * (Actual hours - Standard hours) = 40* (5,000 - 4,750) = 10,000 U